OpenStudy (anonymous):

Ethylene glycol (C2H6O2) is used as an antifreeze in cars. If 400 g of ethylene glycol is added to 4.00 kg of water, what is the molality? Calculate how much the freezing point of water will be lowered. The freezing-point depression constant for water is Kf = –1.86°C/m.

OpenStudy (anonymous):

molar mass of ethylene glycol = 62g mass of ethylene glycol = 400g mass of solvent (in kg) = 4kg molality = 400/(62∗4) =1.61m delta Tf = Kf * m = -1.81 * 1.61 = -2.99 C